Reports of 31355
DGS (General)
satoru.ikeda - 12:06 Friday 23 May 2025 (33880) Print this report
Restart of FEPCs (as a Preventive Measure)

Summary:
 Due to the possibility of system hang-ups after prolonged operation, FEPCs not connected to Dolphin were restarted.
 This is purely a preventive action and does not imply that a hang-up is certain to occur due to long runtimes.
 The FEPCs currently connected to Dolphin are scheduled for restart at a later time.

 Related Entry:
 K-Log #33855: k1px1 was rebooted to recover from a hang-up

Targeted FEPCs:
 k1omc1, k1ix1, k1iy1, k1ex1, k1ey1, k1ex0, k1ey0, k1iy0

FEPCs Excluded from This Restart:
 These systems were recently restarted and therefore were not included this time:
 k1test0: 02 May 2025 (K-Log #33659)
 k1px1: 21 May 2025 (K-Log #33855)
 k1mcf0: 16 April 2025 (K-Log #33438)

Work Details:
 A software reboot was performed using sudo reboot.

Uptime Logs After Restart:
  K1LSC0:  11:08:42 up 371 days, 17:29,  0 users,  load average: 0.15, 0.36, 0.34
  K1ASC0:  11:08:42 up 318 days, 20:18,  0 users,  load average: 0.17, 1.76, 2.18
  K1ALS0:  11:08:42 up 18 days, 21:17,  0 users,  load average: 0.00, 0.03, 0.05
   K1IOO:  11:08:43 up 371 days, 21:45,  0 users,  load average: 0.03, 0.07, 0.06
  K1IOO1:  11:08:43 up 356 days, 20:27,  0 users,  load average: 0.00, 0.01, 0.05
  K1IMC0:  11:08:43 up 371 days, 21:38,  0 users,  load average: 0.00, 0.07, 0.13
   K1PRM:  11:08:43 up 7 days, 18:03,  0 users,  load average: 0.01, 0.02, 0.05
   K1PR0:  11:08:43 up 101 days, 20:59,  0 users,  load average: 0.00, 0.03, 0.05
   K1PR2:  11:08:44 up 371 days, 21:34,  0 users,  load average: 0.07, 0.04, 0.05
    K1BS:  11:08:44 up 384 days, 16:01,  0 users,  load average: 0.00, 0.04, 0.05
   K1SR2:  11:08:44 up 371 days, 22:40,  0 users,  load average: 0.11, 0.08, 0.06
   K1SR3:  11:08:44 up 371 days, 21:51,  0 users,  load average: 0.07, 0.07, 0.05
   K1SRM:  11:08:44 up 371 days, 21:25,  0 users,  load average: 0.05, 0.05, 0.05
  K1OMC0:  11:08:44 up 371 days, 21:21,  0 users,  load average: 1.89, 1.05, 0.65
  K1OMC1:  11:08:45 up 51 min,  0 users,  load average: 0.00, 0.01, 0.05
  K1MCF0:  11:08:45 up 37 days, 46 min,  0 users,  load average: 0.02, 0.02, 0.05
   K1IY0:  11:08:45 up 8 min,  0 users,  load average: 0.01, 0.05, 0.05
   K1IX1:  11:08:45 up 45 min,  0 users,  load average: 0.78, 0.57, 0.40
   K1IY1:  11:08:45 up 40 min,  0 users,  load average: 0.01, 0.20, 0.24
   K1EX0:  11:08:46 up 27 min,  0 users,  load average: 0.03, 0.10, 0.10
   K1EY0:  11:08:46 up 22 min,  0 users,  load average: 0.03, 0.13, 0.10
   K1EX1:  11:08:46 up 35 min,  0 users,  load average: 0.01, 0.18, 0.23
   K1EY1:  11:08:46 up 31 min,  0 users,  load average: 0.01, 0.21, 0.25
   K1PX1:  11:08:46 up 1 day, 18:31,  0 users,  load average: 0.02, 0.04, 0.05
 K1TEST0:  11:08:46 up 20 days, 23:43,  0 users,  load average: 0.02, 0.24, 0.30

 

Non-image files attached to this report
DGS (General)
satoru.ikeda - 11:49 Friday 23 May 2025 (33879) Print this report
Update model files

Request from Ushiba-san
[k1IOO] 
model: k1imcasc
detail: Change DAQ channel item.
 DAQ channel size to be changed:
  Before: 45056 -> After:44032 (-1024)

 [Change]
 REFL_QPDA{1,2}_RF14_{I,Q}_{PIT,YAW} 512->2048
 REFL_QPDA{1,2}_DC_{PIT,YAW}_OUT 256>2048
 MCE_TRANS_QPDA{1,2}_DC_{PIT,YAW}_OUT 256->2048
 IMMT1_TRANS_QPDA{1,2}_DC_SUM_OUT 16384->2048

 [New]
 DOF{1,2,3,4,5}_{P,Y}_OUT 512
 QPDA{1,2}_DC_CTRL_{P,Y}_OUT 512

 [Remove]
 MC{I,E,O}_{PIT,YAQ}_OUT 512
 MCE_TRANS_QPDA{1,2}_DC_SUM_IN1 256
 PZT{1,2}_{PIT,YAW}_OUT 512
 IMMT1_TRANS_QPDA{1,2}_DC_SUM_IN1 256

Request from YamaT-san
[k1TEST0] 
model: k1detdq0
detail: Added one EPICS record as CFC check flag (related to klog#33821)
 

Images attached to this report
Non-image files attached to this report
CAL (General)
hirotaka.yuzurihara - 9:16 Friday 23 May 2025 (33878) Print this report
TCam photo session 20250523

I took the TCam photos for four mirrors at 8:33 ~ 8:42 this morning. The previous work is klog33648.

  • ETMX
  • For these mirrors, I updated the reference positions. I also re-draw the yellow line shown at the k1mon4. For the other mirrors, we don't need to update the reference positions.
  • I didn't know the amount of the MISALIGNED for the ETMX and ETMY suspensions is enough not to reflect the green laser to the input mirror. As I checked the images taken this morning, it's enough to request the MISALIGNED state, not the MISALIGNED_BF state.
CRY (Cryo-payload IX)
shinji.miyoki - 8:49 Friday 23 May 2025 (33877) Print this report
Comment to Stop cryocooler for ITMX payload (33820)

The 1st stage of the REF4/2 cryocooler (4K_REF2/4_50K_HEAD) is now approaching the Tc. So slight H2O release will be expected.

By the way, the recent warm-up speeds of REF2/4 2nd stages (4K_REF2/4_4K_HEAD) are different, while the recent warm-up speeds of the REF2/4 1st stages  (4K_REF2/4_50K_HEAD) are similar. This means that the REF2 2nd stage has more heat invasion than REF4.

Images attached to this comment
AEL (General)
dan.chen - 6:42 Friday 23 May 2025 (33876) Print this report
Comment to Pre-Installation Test Report of High Power Coil Driver for ETMX (33854)

With Shingo Hido

> we will verify the operation of the new circuit.
We measured the TF of the new circuit (S2305285) with changing the relay switchs.
These results look fine.

The file names are showing the measured condition including channel number and filter configurations.
Filter configuration: for instance, "110" means the first and the second filters were ON, but the third one = the last one was OFF, counting from the input side of the circuit.

Images attached to this comment
AEL (General)
dan.chen - 5:30 Friday 23 May 2025 (33875) Print this report
Comment to Pre-Installation Test Report of High Power Coil Driver for ETMX (33854)

Date: May 22, 2025

Summary

Following the pre-installation test of the HPCD (S2315284) for ETMX, we conducted additional investigation to identify the cause of the irregular behavior seen in Channel 3. A flipped 24 kHz input cable was found and fixed, and a malfunction of the relay switch in the final stage of Channel 3 was identified. However, these issues alone do not fully explain the previously observed abnormal transfer function. As a result, we plan to migrate the 24 kHz injection circuit to a new HPCD unit (S2305285), with the filter configuration updated to 3–30 Hz. The migration will begin next week. Before the migration, we will verify the operation of the new circuit.

Details

  • One of the 24 kHz input cables was found to be flipped and has been corrected.
  • The final-stage relay switch in Channel 3 was confirmed to be malfunctioning.
  • However, the irregular transfer function observed previously cannot be fully explained by the switch issue alone.
  • Therefore, we have decided to migrate the 24 kHz injection circuit to another HPCD unit, S2305285.
  • During this migration, the internal filter configuration will also be updated to 3–30 Hz.
  • The migration work is scheduled to begin next week.
  • Prior to that, a functional check of the new HPCD circuit (S2305285) will be conducted.
CAL (General)
takahiro.yamamoto - 22:00 Thursday 22 May 2025 (33874) Print this report
consideration about dewhitening compensation for a new 30Hz/3Hz coil driver

Abstract

We are now planning to use a special coil driver for ETMX which has 30Hz:3Hz zero/pole pairs instead of 10Hz:1Hz pairs.
Pre-test of circuit measurement is ongoing and progress is reported in klog#33854.
We noticed that this special coil driver has different response from normal coil driver not only when dewhitening is enabled but also when it's disabled.
We also found that this response with disabling dewhitening makes ~10% bias as maximum by changing a number of dewhitening stages.
This effect may makes a non-negligible calibration errors in the case of parallel execution of observation and working with observation

It's not so difficult to compensate these two effect only for front-end calibration or only for low-latency calibration.
But an additional consideration about the operation is required to make the correction consistent for both front-end and low-latency calibration.
 

Details

Difference between normal and special coil drivers
zero/pole frequency of dewhitening filter is decided by R7 (should be same value as R25), R11, R19, and C22 in Fig.1.
- zero = 1 / [ 2 * pi * (C22 * 2) * (R7 + (R11 || R19) * 0.5) ]
- pole = 1 / [ 2 * pi * (C22 * 2) * (R11 || R19) * 0.5) ]
For changing zero/pole pairs from 10Hz/1Hz to 30Hz/3Hz, C22 was changed from 4.7uF to 1.5(1.6?)uF.

In this circuit, positive and negative paths are connected via R19 and C22 even when dewhitening is turned off. As a result, frequency dependency appears also in the case of disabling dewhitening as
- zero = 1 / [ 2 * pi * (C22 * 2) * (R7 + R19 * 0.5) ]
- pole = 1 / [ 2 * pi * (C22 * 2) * (R19 * 0.5) ]
These values are ~0.03Hz in the normal coil driver and it can be regarded as flat response above 1Hz. On the other hand, they becomes ~0.1Hz in the special coil driver and it cannot be negligible especially for TF phase. (Response of normal and special coil drivers are shown in Fig.2.) If we will decide to use the special coil driver, we need to compensate them with accurate parameters which will be measured after the installation.

Operation for the parallel execution of observation and working with observation
In the case of normal observation, we can fix a number of used dewhitening. So we can take into account a gain above 10Hz of coil driver without dewhitening filter as a part of the actuator efficiency. But in the case of parallel execution of observation and working with observation, we cannot fix a number of use dewhitening and it might be changed. This effect corresponds to 0.8dB (~10%) as shown in Fig.2. The easiest way to compensate this effect and to keep front-end calibration is applying compensation filters in COILOUTF. But this way seems to be some problem on low-latency calibration according to a chat with Hido-kun. We don't have any good idea yet to keep low-cost to manage both front-end and low-latency calibrations. Though there is not so much time to consider it, we need a better solution.

Images attached to this report
MIF (General)
kenta.tanaka - 18:25 Thursday 22 May 2025 (33873) Print this report
Comment to Initial alignment 20250522 and the interferometer reached the PRFPMI_RF_LOCK (33869)

We suceeded in restoring PRFPMI DC lock with 10W after the ITMX release by warming up.

Fig. 1, 2, 3, and 4 show the OLTFs of DARM, MICH, PRCL, and CARM, respectively. DARM, MICH and CARM seem to be lower to at most ~1dB than previous ones.  PRCL seems to be almost consistent with the previous one.

Therefore, LSCs seems to be fine.

Images attached to this comment
CRY (General)
shinji.miyoki - 17:42 Thursday 22 May 2025 (33872) Print this report
The considearation on how to defrost mirrors

( IX case )

Judging from the IX IM heating-up results with 0.2A after turning off only two cryocoolers for the IX payload, the temperature of IM can be over 160K (> Tc) within 4 days. While, the temperatures of the outer (80 K) shield could be less than Tc at least within 4 days.

On the other hand, in the case of IX, the temperature of the outer (80K) shield is just over the Tc if we operate only two cryocoolers for the outer/inner shields.

So, we expect to be able to defrost mirrors, with the care of the temperature of the outer shields by turning on/off the cryocoolers for the cryopayload to avoid too much H2O evaporation from the outer shield. One concern is the frosting on the widows on the inner shield. Another concern is the too much temperature difference between IM and others.

( IY case )

In the case of IY, the temperature of the outer (80K) shield is just "below" the Tc, even if we operate only two cryocoolers for the outer/inner shields.

So, what we should do is heating up IY IM with 0.2A, after turning off two cryocoolers for the IY cryopayload.  

 

MIF (General)
hirotaka.yuzurihara - 15:37 Thursday 22 May 2025 (33871) Print this report
Comment to Finesse measurement Xarm 250522 (33857)

DATE : 2025/05/22 05:26 UTC
ARM : X
TEMP ITM : 107.4
TEMP ETM : 39.1
NUM : 10
IFO REFL HWP : 152.0
PSL HWP : 173.0
IMC POWER : 6.5
VALUE : 1408.6
ERROR : 8.4

Measured data is stored to /users/Commissioning/data/Finesse/Xarm/20250522-052647


The result is almost consistent with one of this moring.

Images attached to this comment
MIF (General)
hirotaka.yuzurihara - 15:36 Thursday 22 May 2025 (33870) Print this report
Comment to Finesse measurement Yarm 250522 (33860)

DATE : 2025/05/22 06:00 UTC                                                       
ARM : Y
TEMP ITM : 37.5
TEMP ETM : 42.4
NUM : 10
IFO REFL HWP : 152.0
PSL HWP : 173.0
IMC POWER : 6.6
VALUE : 1293.7
ERROR : 8.6

Measured data is stored to /users/Commissioning/data/Finesse/Yarm/20250522-0600374

The result is almost consistent with one of this moring.

Images attached to this comment
MIF (General)
hirotaka.yuzurihara - 13:53 Thursday 22 May 2025 (33869) Print this report
Initial alignment 20250522 and the interferometer reached the PRFPMI_RF_LOCK

[Kenta, Yokozawa, Yuzurihara]
We performed the initial alignment. During taking the alignment for the PRMI, the earthquake happened and the BS was fluctuating ~3urad. We wait for a while and finished the initial alignment at 13:37.
We requested the PRFPMI_RF_LOCK to the LSC_LOCK guardian and confirmed that the interferometer can reach the PRFPMI_RF_LOCK at 13:50.
One concern is that the IR TRANS powers seem to be low <110.

Images attached to this report
Comments to this report:
kenta.tanaka - 18:25 Thursday 22 May 2025 (33873) Print this report

We suceeded in restoring PRFPMI DC lock with 10W after the ITMX release by warming up.

Fig. 1, 2, 3, and 4 show the OLTFs of DARM, MICH, PRCL, and CARM, respectively. DARM, MICH and CARM seem to be lower to at most ~1dB than previous ones.  PRCL seems to be almost consistent with the previous one.

Therefore, LSCs seems to be fine.

Images attached to this comment
PEM (General)
shoichi.oshino - 13:23 Thursday 22 May 2025 (33868) Print this report
Atotsu weather station was stopped from 5/14
Washimi-san found that the Atotsu weather station was stopped from 5/14.
The weather pc did not seem to be turned on, so after turning it on, it came back on successfully.
I checked the logs and found no errors.
VIS (IX)
kenta.tanaka - 13:01 Thursday 22 May 2025 (33867) Print this report
ITMX MN oplev centering and reinsert the resistors between coil drivers and MN coils

Yokozawa, Tanaka

Yokozawa-san performed the centering of MN TILT, LEN, ROLL, and TRA oplevs after the health check.

After that, We re-inserted the resistors between coil drivers and coils of MN. We confirmed that ITMX could reach the ALIGNED state.

VIS (IX)
takafumi.ushiba - 11:05 Thursday 22 May 2025 (33866) Print this report
Comment to Health check ITMX 250522 (33864)

Since suspension would be just released and close to touch somewhere, I tuned the diaggui template for the measurement of pitch TF to reduce the RMS of suspension motion during the TF measurement.
What I did is to add notches at resonances and tuned overall excitation amlitude of the template.
Final setting is shown in fig1.

Figure 2 shows the TF measurement results.
It looks healthy though the coherence at high frequency is not so good due to the reduction of excitation amplitude.
So, the suspension seems to be healthy but still close to rubbing.
I hope the situation will be getting better as increasing the temperature.

Images attached to this comment
VIS (IX)
takaaki.yokozawa - 10:01 Thursday 22 May 2025 (33865) Print this report
Comment to Health check ITMX 250522 (33864)
I increased the excitation count for P measurement, but the situation didn't change.
Images attached to this comment
VIS (IX)
takaaki.yokozawa - 9:10 Thursday 22 May 2025 (33864) Print this report
Health check ITMX 250522
I performed the health check ITMX.

Most part seemed healthy, we need larger amplitude for P measurement?
Images attached to this report
Comments to this report:
takaaki.yokozawa - 10:01 Thursday 22 May 2025 (33865) Print this report
I increased the excitation count for P measurement, but the situation didn't change.
Images attached to this comment
takafumi.ushiba - 11:05 Thursday 22 May 2025 (33866) Print this report

Since suspension would be just released and close to touch somewhere, I tuned the diaggui template for the measurement of pitch TF to reduce the RMS of suspension motion during the TF measurement.
What I did is to add notches at resonances and tuned overall excitation amlitude of the template.
Final setting is shown in fig1.

Figure 2 shows the TF measurement results.
It looks healthy though the coherence at high frequency is not so good due to the reduction of excitation amplitude.
So, the suspension seems to be healthy but still close to rubbing.
I hope the situation will be getting better as increasing the temperature.

Images attached to this comment
VIS (IX)
takaaki.yokozawa - 8:23 Thursday 22 May 2025 (33863) Print this report
IX suspension commissioning 250522
I checked the ITMX status again.

- TM oplev centering by moving mass and BF Yaw setting
By performed it, the set value of BF Yaw became 940, that was very close value before starting the drifting (895)

- Check the OLG MN DAMP P
From the OLG MN DAMP P, it also good suspension.

- Xarm and Yarm alignment.
It was very hard (we didn't perform the alignment for a long time), but we succeeded to perform the initial Xarm and Yarm alignment.

- Finesse measurement
as shown in klog33857, we can lock the IRX and GRX without any problem.

- LOCK_ACQUISITION state
Still we need to perform the MN oplev centering and damping filter tuning, we can set the LOCK_ACQUSITION state to ITMX
Images attached to this report
CRY (General)
shinji.miyoki - 8:19 Thursday 22 May 2025 (33862) Print this report
Comment to All cryostat temperature checks: we should be careful of IX_4K_REF3_50K_HEAD. (33356)

EX_50K_REFBRT_HEAD temperature could not be below 86 K, while 56K in EX_50K_REFARM_HEAD. 

Images attached to this comment
VIS (IX)
shinji.miyoki - 8:15 Thursday 22 May 2025 (33861) Print this report
Comment to Photosensor signals after starting payload cryocooler (33772)

More reports,

  • VIS-ITMX_DAMP_{L}_INMON seemed to reach B
  • VIS-ITMX_DAMP_{R}_INMON seemed to reach A
  • VIS-ITMX_DAMP_{V, P}_INMON seemed to be between A and B.
Images attached to this comment
CRY (Cryo-payload IX)
shinji.miyoki - 8:12 Thursday 22 May 2025 (33859) Print this report
Comment to Stop cryocooler for ITMX payload (33820)

The temps at the 1st heads of the REF2/4 cryocoolers are around 142K (< 157K (Tc)). Not near, but not so far. H2O is now coming from these areas outside the inner shields. So, I hope H2O frosting on mirrors will not be serious.

Images attached to this comment
MIF (General)
takaaki.yokozawa - 8:10 Thursday 22 May 2025 (33860) Print this report
Finesse measurement Yarm 250522
DATE : 2025/05/21 22:42 UTC                                                                                                                                                 
ARM : Y
TEMP ITM : 37.3
TEMP ETM : 42.5
NUM : 10
IFO REFL HWP : 152.0
PSL HWP : 173.0
IMC POWER : 6.3
VALUE : 1290.2
ERROR : 8.2
Measured data is stored to /users/Commissioning/data/Finesse/Yarm/20250521-224243
 
https://gwdet.icrr.u-tokyo.ac.jp/~controls/trendMon/finesse.html
Images attached to this report
Comments to this report:
hirotaka.yuzurihara - 15:36 Thursday 22 May 2025 (33870) Print this report

DATE : 2025/05/22 06:00 UTC                                                       
ARM : Y
TEMP ITM : 37.5
TEMP ETM : 42.4
NUM : 10
IFO REFL HWP : 152.0
PSL HWP : 173.0
IMC POWER : 6.6
VALUE : 1293.7
ERROR : 8.6

Measured data is stored to /users/Commissioning/data/Finesse/Yarm/20250522-0600374

The result is almost consistent with one of this moring.

Images attached to this comment
CRY (Cryo-payload IX)
nobuhiro.kimura - 7:48 Thursday 22 May 2025 (33858) Print this report
Comment to Stop cryocooler for ITMX payload (33820)

[Kimura]

 The increasing trend of H2O in the residual gas component still continues.
For reference, two trend graphs of Q-mass measurements for IXC and BS are attached.
The partial pressure of H2O did not change in the Q-mass on the BS side, suggesting that the H2O component did not flow into the ITMY side due to ITMX heating.

Images attached to this comment
MIF (General)
takaaki.yokozawa - 7:15 Thursday 22 May 2025 (33857) Print this report
Finesse measurement Xarm 250522
Images attached to this report
Comments to this report:
hirotaka.yuzurihara - 15:37 Thursday 22 May 2025 (33871) Print this report

DATE : 2025/05/22 05:26 UTC
ARM : X
TEMP ITM : 107.4
TEMP ETM : 39.1
NUM : 10
IFO REFL HWP : 152.0
PSL HWP : 173.0
IMC POWER : 6.5
VALUE : 1408.6
ERROR : 8.4

Measured data is stored to /users/Commissioning/data/Finesse/Xarm/20250522-052647


The result is almost consistent with one of this moring.

Images attached to this comment
AEL (General)
dan.chen - 7:08 Thursday 22 May 2025 (33856) Print this report
Comment to Pre-Installation Test Report of High Power Coil Driver for ETMX (33854)
Search Help
×

Warning

×